OpenStudy (mwhit2101):

What is 10 to the 320585068702974098930271397948579137498759138751307408943086023348105739874917248450918304983025710980398 power?

OpenStudy (mwhit2101):

Ah I'd really like to know...

OpenStudy (anonymous):

its 10^10^10^2.019 and i'm serious

OpenStudy (mwhit2101):

Can I have it in nums??

OpenStudy (anonymous):

or 3.2x10^104

OpenStudy (anonymous):

like this then? 320 585 068 702 974 098 930 271 397 948 579 137 498 759 138 751 307 408 943 086 023 348 1058 739 874 917 248 450 918 304 983 025 710 980 399

OpenStudy (anonymous):

haha funny question

OpenStudy (mwhit2101):

Hemmmm yes tanks i give u medal

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

hahaha, this thread.... thanks for the laughs

OpenStudy (mwhit2101):

@wio yeah I see how that would be correct rather then his. Wow thats a lot of zeros tahts like a book haha well thanks. :)

OpenStudy (mwhit2101):

@DemolisionWolf You told an untruth bad child!

OpenStudy (anonymous):

Yeah, but you did.

OpenStudy (mwhit2101):

hmmm

thomaster (thomaster):

@wio your reply is so long I can't even delete it xD

OpenStudy (♪chibiterasu):

lol. really? x3

OpenStudy (♪chibiterasu):

~(*O*)~

thomaster (thomaster):

ow woo it's gone now xD finally, took a while to process I guess

OpenStudy (♪chibiterasu):

wio just broke spam levels. Almost indestructable spam :o

OpenStudy (anonymous):

If all my posts were deleted, where do my medals go?

OpenStudy (anonymous):

@wio teach me where i went wrong, i don't want to be the one spreading garbage around on here!

OpenStudy (anonymous):

\[ 10^{320585068702974098930271397948579137498759138751307408943086023348105739874917248450918304983025710980398 } \]Is a number with \[ 320585068702974098930271397948579137498759138751307408943086023348105739874917248450918304983025710980398 \]0s

OpenStudy (anonymous):

And a \(1\) in front.

OpenStudy (anonymous):

\(10^n\) is \(1\) followed by \(n\) \(0\)s.

OpenStudy (anonymous):

\[ 10^0=1\\ 10^1=10\\ 10^2=100\\ 10^3=1000 \]

OpenStudy (anonymous):

awh, i'm with ya!

OpenStudy (anonymous):

Wolfram gave you the number of digits, which is just the original exponent.

OpenStudy (anonymous):

that makes way more sense. in the midst of the excitement i ran w/o looking thanks for the incite!